Calibration update - last line before Run 14 BES

Speaker : Gene V. Buren ( BNL )


Talk time : 12:15, Duration : 00:20
 15 GeV AuAu Calibrations
  • 15 GeV AuAu
    • TPC
      • Drift velocities [done, final]
        • Required manual patching the database for errors in the STAR clock frequency
      • T0 [done, final]
      • Twist (global alignment) [done, final]
        • R&D on alignment would require revisiting this (and all subsequent) calibration(s)
      • SpaceCharge & GridLeak [done, initial]
        • May result in some additional runs not to process (odd RICH scalers); will make public when finalized (this week)
        • Not much lever arm to see east-west asymmetry, but will need a closer look to finalize
      • dE/dx [in progress]
        • R&D on a more accurate fit function to the distribution of cluster-level dE/dx
        • Not a dependency for the preview production
    • EMC
      • ETOW & BTOW pedestals and status tables [done]
        • No emc-check nor L2 sources for these during AuAu15, so we are using values obtained from beginning of AuAu200 operations
    • TOF
      • VPD slewing [done]
        • Needed for sufficiently precise online triggering
      • BTOF alignment [near done]
      • VPD and T0 offsets [ongoing]
      • ...but will use "startless" timing for the preview production
    • HFT
      • Not part of the preview production request
  • 200 GeV AuAu
    • TPC
      • Drift velocities [paused]
        • Calibrated for March 14-23
        • ...but the STAR clock frequency problems make the drift velocity wrong, and its application wrong, so we were populating the database with values that will need replaced
        • Ongoing activity to resolve the problems by using a different frequency source
          • Will patch the existing database values at the same time that we switch to the new source for live feeds
      • T0 [near done, initial]
        • Small shift (~15 ns) expected, and observed
      • Twist [done, initial]
        • Using 15 GeV calibration, as the magnetic field is the same (standard practice)
        • Will re-visit when HFT (support)material accounting in tracking is in place
      • SpaceCharge & GridLeak [done, initial]
        • Using an east-west asymmetric calibration
          • Much larger asymmetry than CuAu! (dAu ~20%, CuAu ~4%, AuAu with HFT ~13%)
        • R&D needed for final calibration
      • dE/dx [awaits final SpaceCharge & GridLeak]
    • TOF
      • Awaits final TPC calibrations relevant for track projections
    • HFT
      • Dead & hot channel masking [ongoing]
      • Alignment work [ongoing, as presented last week]
        • Looking more carefully now at HFT - TPC alignment
    • EMC
      • ETOW & BTOW pedestals and status tables [done, initial]
        • Initial calibration is the same used for 15 GeV, but just one static set
        • Pushing for emc-check runs

-Gene